OpenStudy (lcorazon1024):

Complete the following statement. Write your answer as a decimal or whole number. % of $10 = $9

OpenStudy (madisongoode):

Okay, so, it's pretty simple. (A / 100) x B = C Basic Equation Solve for A : (A / 100) = C / B A = (C / B) x 100 A = (9 / 10) x 100 A = 90%

OpenStudy (madisongoode):

90% = 90/100 = 9/10

OpenStudy (madisongoode):

Hope this helps. And your welcome.